# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v3

@@ IV-C PARC: Avoid Set without Tracking Error
Without accounting for tracking error, the avoid set is the set of pairs of initial planning model state and trajectory parameter such that any plans generated from outside the avoid set are guaranteed to not collide with the obstacles at all continuous time before $t_{\mathrm{\textnormal{f}}}$. Recall that, per (8), we model the continuous-time behavior as straight line segments between the discrete states of the PWA system.
Intuitively, the avoid set identifies all trajectories where the straight lines connecting the PWA planning model’s discrete states collide with any obstacles:
Proposition 11 (Unsafe Plans w/o Tracking Error).
[…]
then
Proof.
By the definition of convex-hull, we have
[…]
$\mathbf{x}(t),\mathbf{x}(t+\Delta t)\in\overline{\Omega}_{t}\bigcup\overline{\Omega}_{t+\Delta t}\},$ (26a)
[…]
Thus (24) necessarily implies (25). ∎
Next, before we can overapproximate the avoid set between two-time steps, we must adapt our notion of extended translation invariance to the PWA planning model:
Lemma 12 (PWA ETI).
@@
$\left\|\hat{\mathbf{C}}_{i,t}^{W}{(\hat{\mathbf{C}}_{i,t}\mathbf{x}+\mathbf{d}_{i,t})_{1:{n}_{w}}}\right\|_{\mathrm{\textnormal{2}}}=0\quad\forall\mathbf{x}\in{\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})}$ (27)
where $\hat{\mathbf{C}}_{i,t}=\mathbf{C}_{i,t}-\mathbf{I}_{{n}_{x}}$ and $\hat{\mathbf{C}}^{W}_{i,t}$ denotes the ${n}_{w}\times{n}_{w}$ upper-left submatrix of $\hat{\mathbf{C}}_{i,t}$, namely $\hat{\mathbf{C}}_{i,t}^{W}\triangleq(\hat{\mathbf{C}}_{i,t})_{1:{n}_{w},1:{n}_{w}}$
@@
The definition of PWA-ETI (27) in other subspace is similarly defined by substituting $W$ to the subspace of interest. Note that $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ is an ETI system in $W\times K$ since the trajectory parameter $\mathbf{k}\in K$ is constant.
Now, for avoid-set computation, it is helpful to relabel the ETI states of $\mathbf{x}(t)$, identified per Lemma 12:
$\mathbf{x}(t)$ $=\begin{bmatrix}\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\\ \mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\end{bmatrix},$ (28)
where $\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\in X_{\mathrm{\textnormal{ETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}$ are the first ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}$ states in $\mathbf{x}(t)$ that are ETI states, and $\mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\in X_{\mathrm{\textnormal{nonETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}$ are the remaining states in $\mathbf{x}(t)$. Thus ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}\geq{n}_{w}+{n}_{k}$. Note that all sets and matrices including $\mathcal{O}$, and $P_{\mathrm{\textnormal{G}}}$ are relabeled accordingly. This reordering is only to simplify the exposition.
We are now ready to state our first key contribution: overapproximating the avoid set without tracking error. The following theorem overapproximates the continuous-time BRS of an obstacle and intersects it with the discrete-time reach-set of the PWA planning model to identify unsafe plans.
@@
In summary, Proposition 11 establishes a necessary condition for the collision to occur between two discrete timesteps, while Theorem 13 overapproximates the avoid sets contained in each intermediate BRS.
@@
We now propose a straightforward scheme to construct the avoid set using Proposition 11 and Theorem 13: reference trajectories generated from outside the avoid set are guaranteed to avoid the obstacle. First, for each timestep $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and for each obstacle $i=1,\cdots,{n}_{\mathcal{O}}$, we check if (25) is true; this reduces conservativeness because we identify obstacles for which we do not have to compute the avoid set $\overline{\Lambda}_{i,t,t}$ with Theorem 13, at the expense of additional convex hull computations. That is, one can speed up computation but potentially make the avoid set more conservative by skipping checking (25). Then, we apply the following corollary:
@@
$\overline{\Lambda}_{i,t,j}$ $=\mathcal{B}\!\left(\overline{\Lambda}_{i,t,j+\Delta t},\mathbf{C}_{s_{j},j},\mathbf{d}_{s_{j},j}\right)\ \forall j=0,\Delta t,\cdots,t-\Delta t.$ (31)
Then, the final, overapproximated avoid set $\overline{\Lambda}$ can be stored as a union of H-polytopes as:
[…]
@@
This follows as a consequence of Theorem 13 and since the BRS operation, $\mathcal{B}$ is exact for an affine system defined for a mode sequence (Lemma 1). ∎
To sample from within the reach set but not from the avoid set, one can either rejection-sample, sample via random-walk algorithms [kannan2009random, chen2018fast], or compute the set difference between the reach set and the avoid set [baotic2009polytopic], and then sample via solving an LP. An example of the reach set and the avoid set are shown in Fig. 3; notice that the avoid set is a nonconvex union of polytopes.
Figure 3: (Left) Visualization of the reach set (green) and the avoid set (red) computation from Section IV-C for the TurtleBot example (Section III-D), sliced in the first three dimensions with respect to the initial condition $\mathbf{p}_{0}=[-3.5,-0.5,\pi/5]^{\intercal}$. The black star denotes the expert plan with $\mathbf{k}=[-0.318,0.9]^{\intercal}$ and the blue star denotes a safe plan with $\mathbf{k}=[-0.320,1.1]^{\intercal}$. (Right) Illustration of the resulting trajectory in $p_{x}$ and $p_{y}$, with the expert plan shown as a dashed black line and the safe plan shown as a dashed blue line. While the expert plan collides with the obstacle (red), PARC is still able to compute a safe plan with the same mode sequence that safely reach the goal set (green).
Figure 4: Visualization of the reach set (green), the avoid set (red), and a grid-based over-approximation of the true avoid set (yellow) for the TurtleBot example (Section III-D) for different values of the timestep $\Delta t$, with the same problem setup and slicing as Fig. 3. The grid-based over-approximation was obtained by dividing the reach set into a $50\times 50$ grid, and checking if the trajectories from each grid collides with the obstacle using FRS computation [herceg2013multi] and Proposition 11. Both the reach set and the numerical overapproximation error gets smaller as $\Delta t$ decreases.
@@ IV-D Estimating Tracking Error
Now that we have a basis for computing the BRAS, we will model tracking error such that, when incorporated into the PARC framework, we can produce plans that, when followed by the tracking model, guaranteed to reach the goal region while avoiding all obstacles.
@@ IV-E1 PARC: Reach Set with Tracking Error
$\mathbf{\psi}\!\left(\mathbf{x}_{0}\right)\in P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}}.$ (39)
@@
Consider a mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$.. Then we can underapproximate the reach set $\Omega_{0}$ corresponding to this mode sequence by:
@@
$\Omega_{t}=$ $\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$
@@ IV-E2 PARC: Avoid Set with Tracking Error
Proposition 19 (Unsafe Plans w/ Tracking Error).
Consider the buffered obstacle $\tilde{\mathcal{O}}_{t,i}$ for some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and the mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS ${\Omega}_{t}$ and the $(t_{\mathrm{\textnormal{f}}}-t-\Delta t)$-time BRS ${\Omega}_{t+\Delta t}$ using (40). If any planning trajectories between ${\Omega}_{t}$ and ${\Omega}_{t+\Delta t}$ collide with $\tilde{\mathcal{O}}_{t,i}$, that is, if $\exists\ \mathbf{x}(t)\in{\Omega}_{t},\mathbf{x}(t+\Delta t)=\mathbf{C}_{s_{t},t}\mathbf{x}(t)+\mathbf{d}_{s_{t},t}\in{\Omega}_{t+\Delta t}$ such that
[…]
then
$\tilde{\mathcal{O}}_{t,i}\cap\mathrm{\textnormal{conv}}\!\left({\Omega}_{t},{\Omega}_{t+\Delta t}\right)\neq\emptyset.$ (43)
Proof.
This follows directly from Proposition 18 and Proposition 11. ∎
Finally, we can state our second key contribution: overapproximating the avoid set with tracking error (i.e., computing the BRAS).
@@
Follows from Proposition 19 and Theorem 13. ∎
@@
With the avoid set and the reach set from Section IV-E1 computed, if the robot tracks any plans starting in $\Omega_{0}\setminus\Lambda$ it is guaranteed to reach the goal region while avoiding all obstacles. Thus, the set $\Omega_{0}\setminus\Lambda$ is the desired object $\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})$. Next, we demonstrate PARC’s BRAS computation on a variety of numerical examples.
V Experiments
We now assess the utility of our proposed PARC method for generating safe trajectory plans near danger. We seek to understand the impact of planning model design (Section V-A), cooperative vs. adversarial planner-trackers (Section V-B), tight approximation of planning model reachable sets (Section V-C), and respect of time-varying tracking error in planning level (Section V-D). We also assess the utility of learning-based reach-avoid in Sections V-A and V-B.
To ensure fair assessment, we apply PARC to robots from the safe motion planning and control literature: a near-hover quadrotor in 2-D [dawson2022safe] and 3-D [chen2021fastrack], and a more general quadrotor model in 3-D [kousik2019safe]. Implementation details are reported in Appendix -E. All experiments, and training of learning-based methods, are run on a desktop computer with a 24 core i9 CPU, 32 GB RAM, and an Nvidia RTX 4090 GPU.
Key result: PARC outperforms other methods in goal reaching and safety in near-danger scenarios, but requires careful choice of planning model. While it can still guarantee reach-avoid in non-near-danger scenarios, its performance is worse than existing methods. Most experiments are summarized in Table II.
V-A Impact of Planning Model Design
This experiment seeks to understand the impact of choosing a planning model for a given tracking model. We deploy PARC on a tracking model and environment that have been successfully solved by a Neural CLBF [dawson2022safe] approach.
Key result: Naïve design of PARC’s planning model can impact performance negatively. We confirm that a better choice of planning model solves the same task with higher performance in Appendix -G.
V-A1 Experiment Setup
To examine the importance of the planning model, we employ the same dynamics for both the planning and tracking model, namely 2-D near-hover quadrotor dynamics:
The workspace states are $\mathbf{w}=\begin{bmatrix}p_{x}&p_{y}\end{bmatrix}^{\intercal}$, and the inputs are left- and right-side propeller forces
$\mathbf{u}=\mathbf{k}$ $=\begin{bmatrix}{F_{\mathrm{\textnormal{l}}}}&{F_{\mathrm{\textnormal{r}}}}\end{bmatrix}^{\intercal},$
which are held constant for $[0,t_{\mathrm{\textnormal{f}}}]$ to generate plans. The feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ produces time-varying left- and right-side propeller forces using an iLQR controller [kong2021ilqr].
[…]
We use the obstacle and goal environment from [dawson2022safe, Sec. 6.3], wherein the quadrotor must navigate over an obstacle and through a small gap to a goal region (see the left side of Fig. 5 and Fig. 6).
We use PARC to compute a BRAS for this scenario with an expert plan, accounting and without accounting for tracking error. The planning model piece-wise affinize (54) with 80 PWA regions, timestep $\Delta t=0.1$units, and final time $t_{\mathrm{\textnormal{f}}}=2$units. We assess the computation time, as well as the size of the BRAS to investigate the impact of planning model design.

### v4

@@ IV-C PARC: Avoid Set without Tracking Error
Without accounting for tracking error, the avoid set is the set of pairs of initial planning model state and trajectory parameter such that any plans generated from outside the avoid set are guaranteed to not collide with the obstacles at all continuous time before $t_{\mathrm{\textnormal{f}}}$.
Before we can overapproximate the avoid set between two-time steps, we must adapt our notion of extended translation invariance to the PWA planning model:
Lemma 11 (PWA-ETI in Workspace).
@@
$\left\|\hat{\mathbf{C}}_{i,t}^{W\times K}{(\hat{\mathbf{C}}_{i,t}\mathbf{x}+\mathbf{d}_{i,t})_{1:({n}_{w}+{n}_{k})}}\right\|_{\mathrm{\textnormal{2}}}=0\quad\forall\mathbf{x}\in X$ (24)
where $\hat{\mathbf{C}}_{i,t}=\mathbf{C}_{i,t}-\mathbf{I}_{{n}_{x}}$ and $\hat{\mathbf{C}}^{W\times K}_{i,t}$ denotes the $({n}_{w}+{n}_{k})\times({n}_{w}+{n}_{k})$ upper-left submatrix of $\hat{\mathbf{C}}_{i,t}$.
@@
To alleviate the conservativeness of our approach, it is beneficial to extend the ETI property to subspaces beyond the workspace $W$. If the planning model is ETI in subspace $X_{\mathrm{\textnormal{ETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}$, it must fulfill the following conditions:
$\left\|\hat{\mathbf{C}}_{i,t}^{X_{\mathrm{\textnormal{ETI}}}}(\hat{\mathbf{C}}_{i,t}\mathbf{x}+\mathbf{d}_{i,t})_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\right\|_{\mathrm{\textnormal{2}}}$ $=0\quad\forall\mathbf{x}\in X$ (25a)
$\exists X_{\mathrm{\textnormal{nonETI}}}\quad\mathrm{\textnormal{s.t.}}\quad X$ $=X_{\mathrm{\textnormal{ETI}}}\times X_{\mathrm{\textnormal{nonETI}}}.$ (25b)
where $\hat{\mathbf{C}}^{X_{\mathrm{\textnormal{ETI}}}}_{i,t}$ denotes the ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}\times{{n}_{x}}_{\mathrm{\textnormal{ETI}}}$ upper-left submatrix of $\hat{\mathbf{C}}_{i,t}$. We refer to the first ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}$ states of $\mathbf{x}\in X$ as ETI states. Note that $\mathbf{w}$ and $\mathbf{k}$ are ETI states because of Lemma 11 and $X=W\times K\times P_{\mathrm{\textnormal{other}}}$ from (18).
To lower PARC’s conservativeness, we reorder $\mathbf{x}(t)$ such that the most ETI states are placed at the beginning of $\mathbf{p}_{\mathrm{\textnormal{other}}}$, as identified by (25). To simplify exposition for avoid-set computation, we now denote $\mathbf{x}(t)$ as:
$\mathbf{x}(t)$ $=\begin{bmatrix}\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\\ \mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\end{bmatrix},$ (26a)
$\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)$ $=\begin{bmatrix}\mathbf{w}(t)\\ \mathbf{k}\\ {\mathbf{p}_{\mathrm{\textnormal{other}}}}_{\mathrm{\textnormal{ETI}}}(t)\end{bmatrix}$ (26b)
where ${\mathbf{p}_{\mathrm{\textnormal{other}}}}_{\mathrm{\textnormal{ETI}}}(t)$ are the first ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}-{n}_{w}-{n}_{k}$ states in $\mathbf{p}_{\mathrm{\textnormal{other}}}(t)$ that are ETI states, and $\mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\in X_{\mathrm{\textnormal{nonETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}$ are the remaining states in $\mathbf{p}_{\mathrm{\textnormal{other}}}(t)$. Thus, ${{n}_{x}}_{\mathrm{\textnormal{nonETI}}}={n}_{x}-{{n}_{x}}_{\mathrm{\textnormal{ETI}}}$. Note that all sets and matrices including $\mathcal{O}$, and $P_{\mathrm{\textnormal{G}}}$ are relabeled accordingly.
We are now ready to state our first key contribution: overapproximating the avoid set without tracking error. The following theorem overapproximates the continuous-time BRS of an obstacle and intersects it with the discrete-time reach-set of the PWA planning model to identify unsafe plans:
@@
In essence, the intermediate avoid set from (27) overapproximates the avoid set contained in each intermediate BRS. Under specific circumstances, the intermediate avoid set can be computed without convex hull or projection. See Appendix -E2 for details.
@@
We now propose a straightforward scheme to construct the avoid set using Theorem 12, where reference trajectories generated from outside the avoid set are guaranteed to avoid the obstacle:
@@
$\overline{\Lambda}_{i,t,j}$ $=\mathcal{B}\!\left(\overline{\Lambda}_{i,t,j+\Delta t},\mathbf{C}_{s_{j},j},\mathbf{d}_{s_{j},j}\right)\cap\mathcal{P}(\mathbf{A}_{s_{j},j},\mathbf{b}_{s_{j},j}),$ (29)
for all $j=0,\Delta t,\cdots,t-\Delta t$. Then, the final, overapproximated avoid set $\overline{\Lambda}$ can be stored as a union of H-polytopes as:
$\overline{\Lambda}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\bigcup_{t=0}^{t_{\mathrm{\textnormal{f}}}-\Delta t}{\overline{\Lambda}_{i,t,0}}.$ (30)
@@
This follows as a consequence of Theorem 12 and the BRS operation in (29), which is exact for an affine system defined for a mode sequence (Lemma 1). ∎
To sample from within the reach set but not from the avoid set, one can either rejection-sample, sample via random-walk algorithms [kannan2009random, chen2018fast], or compute the set difference between the reach set and the avoid set [baotic2009polytopic], and then sample via solving a linear program (LP). An example of the reach set and the avoid set is shown in Fig. 3; notice that the avoid set is a union of polytopes, hence can be non-convex.
Figure 3: (Left) Visualization of the reach set (green) and the avoid set (red) computation from Section IV-C for the TurtleBot example (Section III-D), sliced in the first three dimensions with respect to the initial condition $\mathbf{p}_{0}=[-3.5,-0.5,\pi/5]^{\intercal}$. The black star denotes the expert plan with $\mathbf{k}=[-0.318,0.9]^{\intercal}$ and the blue star denotes a safe plan with $\mathbf{k}=[-0.320,1.1]^{\intercal}$. (Right) Illustration of the resulting trajectory in $p_{x}$ and $p_{y}$, with the expert plan shown as a dashed black line and the safe plan shown as a dashed blue line. While the expert plan collides with the obstacle (red), PARC is still able to compute a safe plan with the same mode sequence that safely reaches the goal set (green).
Figure 4: Visualization of the reach set (green), the avoid set (red), and a grid-based over-approximation of the true avoid set (yellow) for the TurtleBot example (Section III-D) for different values of the timestep $\Delta t$, with the same problem setup and slicing as Fig. 3. The grid-based over-approximation was obtained by a slower method detailed in Appendix -E1 with forward reachable set (FRS) verification on the reach set, divided into a $50\times 50$ grid. Both the reach set and the numerical overapproximation error gets smaller as $\Delta t$ decreases.
@@ IV-D Estimating Tracking Error
Now that we have a basis for computing the BRAS, we will model tracking error such that, when incorporated into the PARC framework, can produce plans that, when followed by the tracking model, guaranteed to reach the goal region while avoiding all obstacles.
@@ IV-E1 PARC: Reach Set with Tracking Error
$\mathbf{x}(t_{\mathrm{\textnormal{f}}};\mathbf{x}_{0})\in P_{\mathrm{\textnormal{G}}}\ominus E_{t_{\mathrm{\textnormal{f}}}}.$ (37)
@@
Consider a mode sequence $S=(s_{0},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Then we can underapproximate the reach set $\Omega_{0}$ corresponding to this mode sequence by:
@@
$\Omega_{t}=$ $\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)\cap\mathcal{P}(\mathbf{A}_{s_{t},t},\mathbf{b}_{s_{t},t})$
@@ IV-E2 PARC: Avoid Set with Tracking Error
We can finally state our second key contribution: overapproximating the avoid set with tracking error (i.e., computing the BRAS):
@@
Follows from Theorem 12. ∎
@@
${\Lambda}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\bigcup_{t=0}^{t_{\mathrm{\textnormal{f}}}-\Delta t}{{\Lambda}_{i,t,0}}.$ (43)
With the avoid set and the reach set from Section IV-E1 computed, if the robot tracks any plans starting in $\Omega_{0}\setminus\Lambda$, it is guaranteed to reach the goal region while avoiding all obstacles. Thus, the set $\Omega_{0}\setminus\Lambda$ is the desired object $\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})$. Next, we demonstrate PARC’s BRAS computation on a variety of numerical examples.
V Experiments
We now assess the utility of our proposed PARC method for generating safe trajectory plans near danger. We seek to understand the impact of planning model design (Section V-A), cooperative vs. adversarial planner-trackers (Section V-B), tight approximation of planning model reachable sets (Section V-C), and respecting time-varying tracking error in planning level (Section V-D). We also assess the utility of a learning-based reach-avoid method in Section V-A and Section V-B.
To ensure fair assessment, we apply PARC to robots from the safe motion planning and control literature: a near-hover quadrotor in 2-D [dawson2022safe] and 3-D [chen2021fastrack], and a more general quadrotor model in 3-D [kousik2019safe]. Implementation details are reported in Appendix -F. All experiments, and training of learning-based methods, were run on a desktop computer with a 24-core i9 CPU, 32 GB RAM, and an Nvidia RTX 4090 GPU.
Key result: PARC outperforms other methods in goal reaching and safety in near-danger scenarios but requires careful choice of planning model. While it can still guarantee reach-avoid in non-near-danger scenarios, its performance can be worse than existing methods. Most experiments are summarized in Table II.
V-A Impact of Planning Model Design
This experiment seeks to understand the impact of choosing a planning model for a given tracking model. We deployed PARC on a tracking model and environment that have been successfully solved by a Neural CLBF [dawson2022safe] approach.
Key result: Naïve design of PARC’s planning model can impact performance negatively. We confirm that a better choice of planning model solves the same task with higher performance in Appendix -H.
V-A1 Experiment Setup
To examine the importance of the planning model, we employed the same dynamics for both the planning and tracking model, namely 2-D near-hover quadrotor dynamics:
The workspace states are $\mathbf{w}=[p_{x},p_{y}]^{\intercal}$, and the inputs are left- and right-side propeller forces
$\mathbf{u}=\mathbf{k}$ $=[{F_{\mathrm{\textnormal{l}}}},{F_{\mathrm{\textnormal{r}}}}]^{\intercal},$
which were held constant for $[0,t_{\mathrm{\textnormal{f}}}]$ to generate plans. The feedback controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}$ produces time-varying left- and right-side propeller forces using an iLQR controller [kong2021ilqr].
@@ V-A1 Experiment Setup
We used the obstacle and goal environment from [dawson2022safe, Sec. 6.3], wherein the quadrotor must navigate over an obstacle and through a small gap to a goal region (see the left side of Fig. 5 and Fig. 6).
We used PARC to compute a BRAS for this scenario with an expert plan, accounting and without accounting for tracking error. The planning model piecewise-affinized (50) with 80 PWA regions, timestep $\Delta t=0.1$units, and final time $t_{\mathrm{\textnormal{f}}}=2$units. We assessed the computation time, as well as the size of the BRAS to investigate the impact of planning model design.